OpenStudy (anonymous):

Explain, in complete sentences, which method you would use to solve the following system of equations and why you chose that method. Provide the solution to the system. x - 3y + 4z = -9 2x + y + 3z = 6 x - 4y + 3z = -9

rishavraj (rishavraj):

use determinant to solve system of linear equation..the Cramer's rule

OpenStudy (anonymous):

ya i've gotten it since then. can you help me with another?

rishavraj (rishavraj):

didnt u get the answer???

OpenStudy (anonymous):

Yeah I did i just need help with another question I have to do

rishavraj (rishavraj):

which one??

OpenStudy (anonymous):

Solve the following system of equations. x + y - 4z = -7 2x + 3y - 3z = -6 2x + 2y - z = 0

OpenStudy (anonymous):

????

rishavraj (rishavraj):

hold on..lemme solve it..

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

here's a list of possible answers to help you out

OpenStudy (anonymous):

any luck?

rishavraj (rishavraj):

got the answer.,...

OpenStudy (anonymous):

i got B

rishavraj (rishavraj):

b (-3,2,2)??? this one is wrong..

OpenStudy (anonymous):

o well i messed up somewhere then /:

rishavraj (rishavraj):

yeah for sure..just recheck it...

OpenStudy (anonymous):

wait, is it D?

rishavraj (rishavraj):

nope its not D

OpenStudy (anonymous):

lemme check again. bare with me

OpenStudy (anonymous):

okok i think i got it

rishavraj (rishavraj):

which option??

OpenStudy (anonymous):

after a few adjustments i got A but i think i might of messed something up in the 2nd equation

rishavraj (rishavraj):

sorry but its not A...

OpenStudy (anonymous):

omg i'm so bad at this

OpenStudy (anonymous):

i'm an english person so math is hard for me

rishavraj (rishavraj):

u need a lot of practice in the same....i think u did calculation mistake...were u calculating by urself or using any gadget??

OpenStudy (anonymous):

by myself

rishavraj (rishavraj):

good then....try to calculate again..and mind the values

OpenStudy (anonymous):

ok i will look over it again, thanks so much ^O^

rishavraj (rishavraj):

anytime!!!